alones (alones):

@agent0smith the hypotenuse of a right triangle

OpenStudy (darkparkai01):

a^2 +b^2 =c^2

OpenStudy (darkparkai01):

C is the hypotenuse

OpenStudy (anonymous):

He hasn't put the question yet...

alones (alones):

Find the length of the hypotenuse of a right triangle whose legs are 3 and \[\sqrt{2}\] .

OpenStudy (agent0smith):

Use what @DarkParkai01 posted. hypotenuse is c.

OpenStudy (anonymous):

Well we need to square each number. First we square 3 and get 9. Then we square the squareroot of 2 to get 2. You add than and you get 11. Then squareroot 11, or just leave it square root of 11.

alones (alones):

OOOooooh i now i ssee thanks